OpenStudy (anonymous):

Simplify the expression. 2^3 • 2^5 A. 2^8 B. 4^15 C. 2^15 D. 4^8 I'm confused, am i supposed to add them or multiply them?

OpenStudy (anonymous):

Hey there! Those look like exponents. So, 2^8 (2 to the eighth power) is 2*2*2*2*2*2*2*2. Basically, the number multiplied by itself the amount of times the exponent is. Does that make sense?

OpenStudy (ranga):

\[\Large a^m \times a^n = a^{m+n}\]

OpenStudy (ranga):

\[\Large 2^3 \times 2^5 = 2^{3 + 5} = 2^8\]

OpenStudy (anonymous):

So it's 2^8 ?

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

Thank you :)

OpenStudy (ranga):

You are welcome.